ganeshie8 (ganeshie8):

the expression ax^3-8x^2+bx+6 is divisible by x^2-2x-3.Find the values of a and b. @eric_d

ganeshie8 (ganeshie8):

say `g(x) = x^2-2x-3` we're given `g(x)` divides the given expression, so can we say that the individual factors of `g(x)` also divide the given expression ?

OpenStudy (ikram002p):

*

OpenStudy (eric_d):

ok..

ganeshie8 (ganeshie8):

can you factor `g(x)` ?

OpenStudy (eric_d):

(x-3)(x+1)

ganeshie8 (ganeshie8):

Excellent ! that means `x-3` and `x+1` are factors of given expression, right ?

OpenStudy (eric_d):

ys

OpenStudy (ikram002p):

i dnt know how u manage that much of question :-\

OpenStudy (eric_d):

i dnt know how u manage that much of question

OpenStudy (eric_d):

You mean Ganesh ? @ikram002p

OpenStudy (ikram002p):

hehe , yeah :P

OpenStudy (eric_d):

:P

ganeshie8 (ganeshie8):

good :) so we appeal to factor/remainder theorem : if `x-a` is a factor of f(x), then `f(a) = 0`

OpenStudy (eric_d):

ok..

OpenStudy (kainui):

Well isn't this like saying if we have the number 72 and I tell you it is divisible by 6 does this mean that 72 is also divisible by 3 and 2?

OpenStudy (ikram002p):

yep the same kai

ganeshie8 (ganeshie8):

exactly ^^ @eric_d thats a very good analogy to see why the factors of `f(x)` must also divide the given expression

ganeshie8 (ganeshie8):

since `x-3` and `x+1` are factors of given expression also, can we use factor/remainder theorem and get two equations in a and b ? can you try ?

OpenStudy (eric_d):

sure

OpenStudy (eric_d):

27a+3b-66 .....1 -a-b+2 ......2 like this

ganeshie8 (ganeshie8):

Perfect !

ganeshie8 (ganeshie8):

two equations and two unknowns, not very exciting, anybody can solve them :)

OpenStudy (eric_d):

When solving simultaneously, I'm getting a=2.5 ??

OpenStudy (ikram002p):

isolved the question in defferent way , but i got 3 mmm which make no sense cuz this way is correct as well

OpenStudy (eric_d):

You can show ur working here

OpenStudy (ikram002p):

is b = 13 ?

OpenStudy (eric_d):

the given answer is b=-5

OpenStudy (ikram002p):

ok this is what i did 

OpenStudy (ikram002p):



ganeshie8 (ganeshie8):

@eric_d there is a small typo in ur equations which ive overlooked earlier : 27a+3b-66 =0.....1 -a-b\(\color{Red}{-}\)2 =0 ......2 solving gives you http://www.wolframalpha.com/input/?i=solve+27a%2B3b-66%3D0%2C-a-b-2%3D0

OpenStudy (ikram002p):

hehe perfect xD

OpenStudy (eric_d):

So, I can't solve the eqn simultaneously

OpenStudy (eric_d):

@ikram002p

OpenStudy (ikram002p):

-a-b=2 27 a +3b =66 mmm xD -3a-3b =6 + 27 a+3b =66 ـــــــــــــــــــــــــــــــــــــــــ 24a=66+6=27 mmm u know find a

OpenStudy (eric_d):

Okay

OpenStudy (eric_d):

Understood.. Thank you @ganeshie8 @ikram002p @Kainui

OpenStudy (ikram002p):

sorry i made to many typo xD 66+6= 72

OpenStudy (eric_d):

It's alrite